# On Probabilistic Pullback Metrics on Latent Hyperbolic Manifolds

### v2

@@ 4 Hyperbolic Kernel Derivatives
with $g(u)\!=\!\left(\frac{2\rho^{2}}{\nu s^{2}}-\frac{1}{s^{2}}-\frac{u\rho}{s^{3}}\right)$ and $h(u)\!=\!\frac{\text{d}}{\text{d}u}g(u)+\frac{g(u)2\rho}{\nu s}$. The helper functions $g(u)$ and $h(u)$ are essential to understand why standard autodiff tools fail to compute the derivatives (29)-(30). For equal inputs $\bm{x}\!=\!\bm{z}$, the inner product $u$ approaches $-1$, while the distance $\rho=\text{d}_{\mathbb{H}^{D_{x}}_{\mathcal{L}}}(\bm{x},\bm{z})$ and variable $s$ converge to 00. This limit is analytically well behaved for the helper functions, e.g., $\lim_{u\rightarrow-1^{-}}g(u)=\frac{2}{\nu}+\frac{1}{3}$. However, autodiff tools fail to compute the kernel derivatives for equal inputs correctly because they do not cancel out 00-approaching terms. For example, we have for $g(u)$,
@@
In this scenario, autodiff divides by 00, leading to undefined values (NaN). In other cases, e.g., when training a GPLVM, derivatives for equal kernel inputs $\bm{x}=\bm{z}$ may not be that relevant because different latent points rarely become that close. However, these derivatives are essential for computing the Jacobian covariance matrix (20). Although symbolic differentiation libraries could provide a solution, they tend to be too slow for practical purposes. Therefore, the next best approach is to compute the necessary derivatives and their analytical limits manually. The latter are then used when the inputs $\bm{x},\bm{z}$ are closer than a predefined threshold.
Finally, we need the third-order kernel derivatives to optimize pullback geodesics as in the 2D case. We denote the second kernel derivative as $\bm{K}=\frac{\partial^{2}}{\partial\bm{z}\partial\bm{x}}k(\bm{x},\bm{z})\in\mathbb{R}^{4\times 4}$. The third-order derivatives can then be obtained by stacking the individual derivatives of each row and column of $\bm{K}$. The derivative of each row $\bm{K}_{i}$ w.r.t. the first input $\bm{x}$ and each column $\bm{K}_{j}$ w.r.t the second input $\bm{z}$ are given by,
$\!\!\dfrac{\partial\bm{K}_{i}}{\partial\bm{x}}\!=\!\Big{(}\!q(\!u\!)\bm{G}^{\mathcal{L}}\bm{z}\bm{x}^{\!\mathsf{T}}\!\bm{G}^{\mathcal{L}}z_{i}\!+\!h(\!u\!)\bm{G}^{\mathcal{L}}z_{i}\!+\!h(\!u\!)\bm{G}^{\mathcal{L}}\bm{z}\bm{e}_{i}^{\!\mathsf{T}}\!\Big{)}\,\!\Delta_{1i}\!\,e^{\frac{-\rho^{2}}{\nu}}\!,$
$\!\!\dfrac{\partial\bm{K}_{\!j}}{\partial\bm{z}}\!=\!\Big{(}\!q(\!u\!)\bm{G}^{\mathcal{L}}\bm{z}\bm{x}^{\!\mathsf{T}}\!\bm{G}^{\mathcal{L}}x_{\!j}\!+\!h(\!u\!)\bm{G}^{\mathcal{L}}x_{\!j}\!+\!h(\!u\!)\bm{e}_{\!j}\bm{x}^{\!\mathsf{T}}\!\bm{G}^{\mathcal{L}}\!\Big{)}\Delta_{1j}e^{\frac{-\rho^{2}}{\nu}}\!,$
where $q(u)=\frac{\text{d}}{\text{d}u}h(u)+\frac{h(u)2\rho}{\nu s}$ and $\Delta_{ij}$ represents a modified Kronecker delta equal to $-1$ for equal indices and to $1$ otherwise. Stacking the row and column matrices $\bm{K}_{i}$ and $\bm{K}_{j}$ along the first and second dimension accordingly gives the third kernel derivatives $\frac{\partial}{\partial\bm{x}}\bm{K},\frac{\partial}{\partial\bm{z}}\bm{K}\in\mathbb{R}^{4\times 4\times 4}$. Complete derivations are provided in App. B.
Figure 3: Left: Latent embeddings in $\mathbb{H}^{2}_{\mathcal{P}}$ of a subset of the MNIST dataset with $\mathsf{0}$ ( ), $\mathsf{1}$ ( ), $\mathsf{2}$ ( ), $\mathsf{3}$ ( ), $\mathsf{6}$ ( ), and $\mathsf{9}$ ( ). The hyperbolic ( ) and pullback ( ) geodesics interpolate between a $\mathsf{3}$ and a $\mathsf{6}$. Middle: Zoomed-in view to better visualize the metric volume. Right: Ten samples along the decoded geodesics in the image space, corresponding to the hyperbolic (top) and pullback (bottom) geodesic.
@@ 5 Experiments
We test the proposed pullback metric in four distinct experiments to demonstrate that (1) the hyperbolic pullback metric augments the hyperbolic metric with the distortions introduced by the GPHLVM’s nonlinear mapping; and (2) the hyperbolic pullback geodesics adhere to the data distribution, leading to low uncertainty model predictions. Additional experimental details are provided in App. D.
5.1 C-shape
@@
The first experiment serves as a proof of concept to visualize and compare the Euclidean pullback metric $\bm{G}^{\text{P},\mathbb{R}}$ from a GPLVM and the hyperbolic pullback metric $\bm{G}^{\text{P},\mathcal{L}}$ from a GPHLVM. We use a dataset of $N$ 2D $\mathsf{C}$-shape points as both latent variables and observations. In the hyperbolic case, the latent variables and observations are encoded in the Lorentz model, i.e, $\bm{x},\bm{y}\in\mathbb{H}^{2}_{\mathcal{L}}$, but visualized in the Poincaré model which provides a more intuitive representation. We fully specify the LVMs by setting the variance, length scale, and noise variance as $(\tau,\kappa,\sigma_{y}^{2})=(0.7,0.15,0.69)$.
@@
Fig. 2 displays the Euclidean and hyperbolic pullback metric volumes $\sqrt{\text{det}(\bm{G}^{\text{P},\mathbb{R}})}$ and $\sqrt{\text{det}(\bm{G}^{\text{P},\mathcal{L}})}$. In the hyperbolic case, the pullback metric is a $3\times 3$ matrix lying in the 2D tangent space $\mathcal{T}_{\bm{x}^{*}}\mathbb{H}^{2}_{\mathcal{L}}$. Therefore, one of its eigenvalues is always zero, which we exclude to effectively visualize the volume. As shown in Fig. 2-left, the Euclidean pullback metric volume is small near the C-shape data and increases away from it, until a maximum value is reached. The hyperbolic pullback metric volume is also low nearby the data and additionally follows the hyperbolic geometry: It is low near the origin and increases when moving outwards until it becomes infinite at the boundary of the unit circle (see Fig. 2-middle). Overall, the hyperbolic pullback metric effectively integrates the properties of the hyperbolic manifold with those of the data manifold.
@@
Fig. 2 also depicts geodesics generated with and without the pullback metric on both latent manifolds. In both cases, the pullback geodesic closely adheres to the training data. However, in the hyperbolic case, crossing the center experiences a much lower penalty, in terms of curve energy, than in the Euclidean case. Fig. 2-right display the curve energy (23) along the geodesics. We observe that the base manifold geodesics and the pullback geodesics both exhibit constant energy, confirming that the optimized curves are true geodesics on their respective manifolds. For comparison, we also show the curve energy of the base manifold geodesic, but evaluated using the pullback metric tensor. Its varying energy indicates that the base manifold geodesic is not a true geodesic with respect to the pullback metric.
5.2 MNIST Digits
The natural clustering of MNIST images can be viewed as a hierarchy whose nodes are each of the $10$ digit classes. Following this point of view, Mathieu et al. (2019) showed that hyperbolic LVMs are better suited than their Euclidean counterparts for embedding MNIST images. Similarly to previous works (Arvanitidis et al., 2018; Jørgensen & Hauberg, 2021; Lalchand et al., 2022), we here explore the interpolation of handwritten digits from a subset of the MNIST dataset. We embedded the vectorized $28\times 28$ images in $\mathbb{H}^{2}_{\mathcal{L}}$ via a GPHLVM and then use the induced pullback metric.
Fig. 3 shows the learned hyperbolic latent space along with two geodesics interpolating between a digit $\mathsf{3}$ and a digit $\mathsf{6}$. The decoded hyperbolic geodesic and pullback geodesic, sampled at $10$ time steps, show the result of geodesic interpolation in the image space. As expected, the GPHLVM predictions collapse to the non-informative GP mean in regions with sparse data, resulting in blurry images. This is particularly evident along the hyperbolic geodesic, in the first half of the trajectory $(t=[0.11,0.44])$. In contrast, the pullback geodesic tends to avoid empty regions by bending towards the red cluster (digit $\mathsf{9}$), leading to less blurry predictions overall. This shows that the pullback metric better captures the data manifold geometry, and thus may be used, e.g., to measure manifold-aware latent distances and probabilities.
5.3 Multi-cellular Robot Design
Figure 4: Top: Embeddings of multicellular robots from coarse (darker tone) to fine (lighter tone) in $\mathbb{H}^{2}_{\mathcal{P}}$. The embeddings form two clusters originating from an all-vertically-actuated robot ( ) and an all-horizontally-actuated robot ( ). Bottom: Ten samples along the decoded hyperbolic ( ) and pullback ( ) geodesic.
We build the coarse-to-fine framework for designing multicellular robots using hyperbolic embeddings introduced by Dong et al. (2024). Each robot is composed of a $5\times 5$ grid of cells, where each cell can be horizontally actuated (orange), vertically actuated (blue), rigid (black), soft (gray), or empty (white). The design process begins with a fully horizontally or vertically actuated robot and incrementally introduces changes, forming a hierarchical structure that refines from coarse to fine architectures. This hierarchy can then be queried to find suitable robots for specific tasks. Dong et al. (2024) showed that their multicellular robot design approach in hyperbolic space outperformed its Euclidean counterpart in most of the considered design tasks.
The main goal of this experiment is to leverage pullback geodesics as a data augmentation mechanism, so that we can design new robot architectures by decoding geodesics that follow the pattern of previously-designed robots. To do so, we compute the pullback metric from the hierarchical multicellular robot embeddings and optimize geodesics to interpolate between existing robot designs, generating novel ones. Fig. 4 shows the latent space and the mean predictions along both the hyperbolic and pullback geodesics. Note that, while the hyperbolic geodesic produces invalid robot designs, e.g., separated sections at $t=0.22$, the pullback geodesic avoids this issue by complying with the geometry of the data support. This experiment suggests that pullback geodesics can assist the robot design process.
5.4 Hand Grasps Generation
@@
(b) GPHLVM, $\mathbb{H}^{3}_{\mathcal{L}}$
@@
Figure 5: Hand grasps transition from a ring ( ) to a spherical grasp ( ) using hyperbolic ( ) and pullback ( ) geodesics in (a) $\mathbb{H}^{2}_{\mathcal{L}}$ and (b) $\mathbb{H}^{3}_{\mathcal{L}}$. Top left: Latent embeddings. The latent trajectory points are colored according to their corresponding grasp class. The background color in $\mathbb{H}^{2}_{\mathcal{P}}$ represents the pullback metric volume. Top right: Time-series plots of $2$ dimensions of the joint space. The top and bottom rows show the mean of the decoded hyperbolic ( ) and pullback geodesics ( ) with their uncertainty as a gray envelope. A training trajectory to the spherical grasp ( ) and a reversed training trajectory from the ring grasp ( ) are included for reference. Bottom: Generated hand trajectories from the decoded hyperbolic (top) and pullback (bottom) geodesics.
@@
Jaquier et al. (2024) showed that GPHLVM outperform Euclidean GPLVM to learn hand grasp embeddings that comply with the hierarchical structure of human motion taxonomies. Here, we build on this model and explore the use of pullback geodesics as a motion generation mechanism with the aim of creating new motions that transition from one grasp to another. We consider the hand grasp taxonomy of Stival et al. (2019), which organizes common grasps types within a hierarchical taxonomy tree. We use a dataset from the KIT whole-body motion database (Mandery et al., 2016) consisting of $38$ motions of $19$ grasp types obtained from recordings of humans grasping different objects. Each motion corresponds to a subject reaching out to grasp an object from an initial resting pose and results in a trajectory $\bm{Y}\in\mathbb{R}^{N\times 24}$ representing the temporal evolution of the $24$ degrees of freedom of the wrist and fingers.
@@
We train two GPHLVMs augmented with a taxonomy graph-distance prior (Jaquier et al., 2024) and a dynamic prior akin to (Wang et al., 2008) (see App. D.4). Fig. 5 shows the resulting 2D and 3D latent spaces. All training trajectories start near the latent space origin, corresponding to the hand’s initial resting pose, and progress outward until the final grasp. We evaluate the embedding quality via the stress loss $\ell_{\text{stress}}$, that quantifies the preservation of the taxonomy graph distances by the embeddings (see App. D.4 for details). Next, we compare the GPHLVMs with GPLVMs (see also App. D.5). Table 2 shows that the hyperbolic models better capture the taxonomy structure than their Euclidean counterparts, consistent with (Jaquier et al., 2024).
@@
Table 2: Average stress and prediction uncertainty (along decoded geodesics). The stress is compared across geometries for each latent space dimension. The prediction uncertainty is compared between the base manifold and pullback geodesic for each model.
@@
Model Stress $\downarrow$ Prediction Uncertainty $\downarrow$
@@
GPLVM $\mathbb{R}^{2}$ $0.18\pm 0.47$ $0.0524\pm 0.0199$ $\bm{0.0004\pm 0.0002}$
GPHLVM $\mathbb{H}^{2}_{\mathcal{L}}$ $\bm{0.10\pm 0.26}$ $0.0150\pm 0.0227$ $\bm{0.0004\pm 0.0002}$
GPLVM $\mathbb{R}^{3}$ $0.11\pm 0.21$ $0.0030\pm 0.0036$ $\bm{0.0001\pm 0.0001}$
GPHLVM $\mathbb{H}^{3}_{\mathcal{L}}$ $\bm{0.07\pm 0.10}$ $0.0129\pm 0.0150$ $\bm{0.0003\pm 0.0003}$
Fig. 5 shows the hyperbolic and pullback geodesics interpolating from a ring to a spherical grasp, alongside the decoded hand motions. While the hyperbolic geodesic crosses empty regions in the latent space, the pullback geodesic adheres closely to the data support. As both geodesics start and end with the same grasps, their initial and final predictions match. However, the decoded hyperbolic geodesic shows high uncertainty in the middle part of the motion (see Figs. 5(a)-5(b) top right). This is due to the GP reverting to its non-informative mean in regions with no data, where the GP prior dominates the posterior, leading to maximum uncertainty. Consequently, the obtained hand motions display large deviations from the initial ring and final spherical grasps. In contrast, as shown in Fig. 5 and Table 2, the decoded pullback geodesic shows much lower uncertainty, as the geodesic stays closer to the data support. Note that this mirrors the behavior of the GPLVM Euclidean pullback geodesics compared to straight lines (see Fig. 8 in App. D.5 and Table 2), thus highlighting the successful adaptation of the pullback approach to hyperbolic spaces. Moreover, the hand motions obtained from the decoded pullback geodesics result in more realistic transitions that those obtained with the hyperbolic ones. Despite the low uncertainty, the predicted motions sometimes lack smoothness due to the fact that nearby latent points do not always correspond to equally similar hand configurations (i.e., similar joint values). This is mostly alleviated in the 3D latent space, which provides more room to accommodate the latent embeddings.

### v3

@@ 4 Hyperbolic Kernel Derivatives
3D Hyperbolic SE Kernel Derivatives: The 3D hyperbolic SE kernel is given by (3). As in the 2D case, to obtain the expected pullback metric tensor (16), we compute the first two kernel derivatives $\frac{\partial k(\bm{x},\bm{z})}{\partial\bm{x}}|_{\bm{x}=\bm{x}^{*}}$ and $\frac{\partial^{2}k(\bm{x},\bm{z})}{\partial\bm{z}\partial\bm{x}}|_{\bm{x}=\bm{z}=\bm{x}^{*}}$. As derived in App. B.2, these derivatives depend on the function $g(u)\!=\!\left(\frac{2\rho^{2}}{\nu s^{2}}-\frac{1}{s^{2}}-\frac{u\rho}{s^{3}}\right)$ with $\rho=\text{d}_{\mathbb{H}^{D_{x}}_{\mathcal{L}}}(\bm{x},\bm{z})$, $u=\langle\bm{x},\bm{z}\rangle_{\mathcal{L}}$, and $s=\sqrt{u^{2}-1}$. The function $g(u)$ is essential to understand why standard automatic differentiation tools fail to compute the 3D kernel derivatives. For equal inputs $\bm{x}\!=\!\bm{z}$, the inner product $u$ approaches $-1$, while the distance $\rho=\text{d}_{\mathbb{H}^{D_{x}}_{\mathcal{L}}}(\bm{x},\bm{z})$ and variable $s$ converge to 00. This limit is analytically well behaved, i.e., $\lim_{u\rightarrow-1^{-}}g(u)=\frac{2}{\nu}+\frac{1}{3}$. However, automatic differentiation tools fail to compute the kernel derivatives for equal inputs as they do not cancel out 00-approaching terms. For example, we have,
$g(u)=\dfrac{2\rho^{2}}{\nu s^{2}}-\dfrac{1}{s^{2}}-\dfrac{u\rho}{s^{3}}\rightarrow^{\text{ autodiff}}_{\>\>\>\>\bm{x}=\bm{z}}\>\rightarrow\dfrac{0}{0}-\dfrac{1}{0}-\dfrac{0}{0}=\text{NaN}\,,$ (19)
@@
where divisions by 00 lead to undefined values (NaN). While derivatives for equal kernel inputs $\bm{x}=\bm{z}$ may be less relevant when training a GPLVM as different latent points rarely become so close, they are essential for computing the Jacobian covariance matrix $\bm{\Sigma}_{J}$ in (15). Although symbolic differentiation libraries could provide a solution, they tend to be too slow for practical purposes. Instead, we address this issue by computing the necessary derivatives and their analytical limits manually when the inputs $\bm{x},\bm{z}$ are closer than a predefined threshold. Complete derivations are provided in App. B.2.
@@ 5 Experiments
We test the proposed pullback metric in four distinct experiments to demonstrate that (1) the hyperbolic pullback metric augments the hyperbolic metric with the distortions introduced by the GPHLVM’s nonlinear mapping; and (2) the hyperbolic pullback geodesics adhere to the data distribution, leading to low uncertainty model predictions. Note that a direct comparison against other state-of-the-art hyperbolic LVMs [37, 9] is not feasible at this time, as pullback metrics were not derived yet for these models. Additional experimental details are provided in App. D.
@@
C-shape: The first experiment serves as a proof of concept to visualize and compare the Euclidean pullback metric $\bm{G}^{\text{P},\mathbb{R}}$ from a GPLVM and the hyperbolic pullback metric $\bm{G}^{\text{P},\mathcal{L}}$ from a GPHLVM. We design a dataset of 2D $\mathsf{C}$-shape points as both latent variables and observations. In the hyperbolic case, the latent variables and observations are encoded in the Lorentz model, i.e, $\bm{x},\bm{y}\in\mathbb{H}^{2}_{\mathcal{L}}$, but visualized in the Poincaré model for a more intuitive understanding. We fully specify the LVMs by setting the variance, length scale, and noise variance as $(\tau,\kappa,\sigma_{y}^{2})=(0.7,0.15,0.69)$.
@@
Fig. 2 displays the Euclidean and hyperbolic pullback metric volumes $\sqrt{\text{det}(\bm{G}^{\text{P},\mathbb{R}})}$ and $\sqrt{\text{det}(\bm{G}^{\text{P},\mathcal{L}})}$. In the hyperbolic case, the pullback metric is a $3\times 3$ matrix lying in the 2D tangent space $\mathcal{T}_{\bm{x}^{*}}\mathbb{H}^{2}_{\mathcal{L}}$. Therefore, one of its eigenvalues is always zero, which we exclude to effectively visualize the volume. As shown in Fig. 2-left, the Euclidean pullback metric volume is small near the C-shape data and increases away from it, until a maximum value is reached. The hyperbolic pullback metric volume is also low nearby the data and additionally follows the hyperbolic geometry: It is low near the origin and increases when moving outwards until it becomes infinite at the boundary of the unit circle (see Fig. 2-middle). Overall, the hyperbolic pullback metric effectively integrates the properties of both the hyperbolic manifold and the data manifold.
@@
Fig. 2 also depicts geodesics generated with and without the pullback metric on both latent manifolds and the curve energy (17) along the geodesics. In both cases, the pullback geodesic closely adheres to the training data, thus highlighting the successful adaptation of the pullback approach to hyperbolic spaces. However, a key difference lies in the curve energy cost of traversing the center, which is substantially lower in the hyperbolic space. Note that both the base manifold geodesics and the pullback geodesics exhibit constant energy, confirming that the optimized curves are true geodesics on their respective manifolds. For comparison, we also show the curve energy of the base manifold geodesic, but evaluated using the pullback metric tensor. The non-constant energy indicates that the base manifold geodesic is not a true geodesic under the pullback metric.
@@
(b) GPLVM, $\mathbb{R}^{2}$
Figure 3: Left: Embeddings of a subset of the MNIST dataset with digits $\mathsf{0}$ ( ), $\mathsf{1}$ ( ), $\mathsf{2}$ ( ), $\mathsf{3}$ ( ), $\mathsf{6}$ ( ), and $\mathsf{9}$ ( ). The background color represents the pullback metric volume. The base manifold ( ) and pullback ( ) geodesics interpolate between a $\mathsf{3}$ and a $\mathsf{6}$. Right: Ten samples along the decoded geodesics in image space.
MNIST Digits: The natural clustering of MNIST images can be viewed as a hierarchy whose nodes are each of the $10$ digit classes. Following this point of view, Mathieu et al. [37] showed that hyperbolic LVMs are better suited than their Euclidean counterparts for embedding MNIST images. Similarly to previous works [3, 27, 32], we here explore the interpolation of handwritten digits from a subset of the MNIST dataset. We embed $600$ vectorized $28\!\times\!28$ images of $6$ classes into 2D hyperbolic and Euclidean latent spaces using GPHLVM and GPLVM, and then use the induced pullback metrics.
Fig. 3 shows the learned latent spaces along with a base manifold and a pullback geodesic interpolating between a digit $\mathsf{3}$ and a digit $\mathsf{6}$. The decoded geodesics show the result of the interpolation in the image space. As shown in Fig. 3(a), the GPHLVM embeddings tend to be grouped by digits. As expected, the GPHLVM predictions collapse to the non-informative GP mean in regions with sparse data, resulting in blurry images. This is particularly evident along the hyperbolic geodesic for $t=[0.11,0.44]$. In contrast, the hyperbolic pullback geodesic tends to avoid sparse data regions by bending towards the red cluster (digit $\mathsf{9}$), leading to less blurry predictions overall. As shown in Fig. 3(b), the GPLVM embeddings are not as clearly organized as in the hyperbolic latent space and form several groups per digit. As such, despite that the Euclidean pullback geodesic also adheres to the data distribution, the resulting interpolation oscillates between digits. A quantitative comparison of the GPHLVM and GPLVM likelihood values confirms that a hyperbolic geometry is better suited for this dataset (see Table 2). Moreover, the prediction uncertainty values confirm that pullback metrics better captures the data manifold geometry for both models. This validates the effectiveness of hyperbolic pullback geodesics, which thus may be used, e.g., to measure manifold-aware latent distances and probabilities.
Multi–cellular Robot Design:
(a) GPHLVM, $\mathbb{H}^{2}_{\mathcal{L}}$
(b) GPLVM, $\mathbb{R}^{2}$
Figure 4: Left: Embeddings of multi-cellular robots from coarse (darker tone) to fine (lighter tone). The embeddings form two clusters originating from an all-vertically-actuated robot ( ) and an all-horizontally-actuated robot ( ). Right: Ten samples along the decoded base manifold ( ) and pullback ( ) geodesics.
We build the coarse–to–fine framework for designing multi–cellular robots using hyperbolic embeddings as introduced by Dong et al. [16]. Each robot is composed of a $5\times 5$ grid of cells, where each cell can be horizontally actuated (orange), vertically actuated (blue), rigid (black), soft (gray), or empty (white). The design process begins with a fully horizontally or vertically actuated robot and incrementally introduces changes, forming a hierarchical structure that refines from coarse to fine architectures. This hierarchy can then be queried to find suitable robots for specific tasks. Dong et al. [16] showed that their multi-cellular robot design approach in hyperbolic space outperformed its Euclidean counterpart in most of the considered design tasks.
The main goal of this experiment is to leverage pullback geodesics as a data augmentation mechanism, so that we can design new robot architectures by decoding geodesics that follow the pattern of previously–designed robots. To do so, we embed existing robot designs into 2D latent spaces using GPHLVM and GPLVM. These models are trained with a taxonomy graph-distance prior in the form of a stress loss as in [25], to preserve the hierarchical structure of the designs in the latent space (see App. D.3 for details). Then, we compute the pullback metric and optimize geodesics to interpolate between existing robot designs, generating novel ones. Fig. 4 shows the latent space and the mean predictions along both the base manifold and pullback geodesics for each model. As shown in Table 2, GPHLVM achieves a lower stress than GPLVM, indicating that the hyperbolic embeddings better preserve the original hierarchy. For both models, the base manifold geodesics cross sparse data regions, while the pullback geodesics adhere to the data support resulting in low-uncertainty predictions (see Fig. 4 and Table 2). The superior hierarchy preservation in hyperbolic space yields improved transitions between designs, with robot shapes and cell types evolving smoothly along the decoded geodesic (see Fig. 4(a)). In contrast, the decoded Euclidean geodesic leads to abrupt design changes (see Fig. 4(b)). This suggests that hyperbolic pullback geodesics can assist robot design processes.
Hand Grasps Generation:
Refer to caption Refer to caption Refer to caption
(a) GPHLVM, $\mathbb{H}^{2}_{\mathcal{L}}$
Refer to caption Refer to caption Refer to caption
(b) GPLVM, $\mathbb{R}^{2}$
@@
Figure 5: Top left: Embeddings of hand grasps colored according to their corresponding grasp class. The background color represents the pullback metric volume. The base manifold ( ) and pullback ( ) geodesics correspond to a transition from a ring ( ) to a spherical grasp ( ). Top right: Time-series plots of $2$ dimensions of the joint space showing the mean of the decoded geodesics with their uncertainty as a gray envelope. A training trajectory to the spherical grasp ( ) and a reversed training trajectory from the ring grasp ( ) are included for reference. Bottom: Generated hand trajectories from the decoded geodesics.
@@
Jaquier et al. [25] showed that GPHLVMs outperform GPLVMs to learn hand grasp embeddings that comply with the hierarchical structure of human motion taxonomies. Here, we build on this model and explore the use of pullback geodesics as a motion generation mechanism aimed at creating new motions that transition from one grasp to another. We consider the hand grasp taxonomy of [50], which organizes common grasps types within a hierarchical taxonomy tree. We use a dataset from the KIT whole-body motion database [36] consisting of $38$ motions of $19$ grasp types obtained from recordings of humans grasping different objects. Each motion corresponds to a subject reaching out to grasp an object from an initial resting pose and results in a trajectory $\bm{Y}\in\mathbb{R}^{N\times 24}$ representing the temporal evolution of the $24$ degrees of freedom of the wrist and fingers.
@@
We train GPHLVMs and GPLVMs augmented with a taxonomy graph-distance prior [25] and a dynamic prior akin to [52] (see App. D.4 for details). Fig. 5 shows the resulting 2D hyperbolic and Euclidean latent spaces. Fig. 7 in App. D.5 depicts the 3D latent spaces. All training trajectories start near the latent space origin, corresponding to the hand’s initial resting pose, and progress outward until the final grasp. Table 2 shows that the hyperbolic models better capture the taxonomy structure than their Euclidean counterparts, consistent with [25]. Notice that the relatively high stress variance is due to the fact that the taxonomy graph distances of some grasps classes are better preserved by the embeddings than others. However, the distances are better preserved for all classes in hyperbolic latent spaces compared to Euclidean ones (see also App. D.5 and Fig. 8).
@@
Table 2: Marginal likelihood (MNIST), average stress (multi-cellular robots, hand grasps) and prediction uncertainty along decoded geodesics (all). The likelihood and the stress are compared across geometries. The prediction uncertainty is compared between the base manifold and pullback geodesic for each model.
@@
Experiment Model Likelihood $\uparrow$ / Stress $\downarrow$ Prediction Uncertainty (×$\times 100$) $\downarrow$
@@
MNIST Digits GPLVM $\mathbb{R}^{2}$ ${\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}-226.23}$ $9.10\pm 4.28$ $\underline{7.95\pm 3.47}$
GPHLVM $\mathbb{H}^{2}_{\mathcal{L}}$ ${\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\bm{-207.14}}$ $7.31\pm 4.74$ $\bm{5.14\pm 2.21}$
Multi-cellular Robots GPLVM $\mathbb{R}^{2}$ $0.73\pm 1.05$ $3.16\pm 2.77$ $\underline{2.55\pm 2.02}$
GPHLVM $\mathbb{H}^{2}_{\mathcal{L}}$ $\bm{0.58\pm 0.88}$ $4.14\pm 4.65$ $\bm{0.97\pm 1.80}$
Hand Grasps GPLVM $\mathbb{R}^{2}$ $0.18\pm 0.47$ $5.24\pm 1.99$ $\bm{0.04\pm 0.02}$
GPHLVM $\mathbb{H}^{2}_{\mathcal{L}}$ $\bm{0.10\pm 0.26}$ $1.50\pm 2.27$ $\bm{0.04\pm 0.02}$
GPLVM $\mathbb{R}^{3}$ $0.11\pm 0.21$ $0.30\pm 0.36$ $\bm{0.01\pm 0.01}$
GPHLVM $\mathbb{H}^{3}_{\mathcal{L}}$ $\bm{0.07\pm 0.10}$ $1.29\pm 1.50$ $\underline{0.03\pm 0.03}$
Fig. 5 shows the base manifold and pullback geodesics interpolating from a ring to a spherical grasp, alongside the decoded hand motions. While the base manifold geodesic crosses empty regions in the latent space, the pullback geodesic adheres closely to the data support. As both geodesics start and end with the same grasps, their initial and final predictions match. However, the decoded base manifold geodesic shows high uncertainty in the middle part of the motion (see Figs. 5(a)-5(b) top right). This is due to the GP reverting to its non-informative mean in regions with sparse data, where the GP prior dominates the posterior, leading to maximum uncertainty. Consequently, the obtained hand motions display large deviations from the initial ring and final spherical grasps. In contrast, as shown in Fig. 5 and Table 2, the decoded pullback geodesic shows a much lower uncertainty, as the geodesic stays closer to the data support. Moreover, the hand motions obtained from the decoded pullback geodesics produce more realistic transitions than those from the base manifold geodesics, with the hyperbolic pullback geodesics arguably exhibiting the smallest deviations from the start and end grasps. Similar results are obtained with 3D latent spaces, see App. D.5. Despite the low uncertainty, the motions predicted from 2D latent spaces sometimes lack smoothness due to the fact that nearby latent points do not always correspond to equally similar hand configurations. This is alleviated in the 3D latent space, which provides more volume to better accommodate the latent embeddings.